Determine if (-5/2,1) is a solution to -2x+3y=2
Answer by jim_thompson5910(35256)

-2x+3y=2

-2(-5/2)+3(1)=2

10/2+3=2

5+3=2

8=2

Since the equation is false, (-5/2,1) is NOT a solution